Calculate basic and the diluted earnings per common share for the year ended December 31, 2021

. On January 1, 2021, 660 million common shares 20 million 6% cumulative preferred share, $50 par,

6.4 % convertible bond, 3500 million face amount. convertible into 70 million common shares.

The following additional information is available on September 1, 2021,

Canaday sold 72 million additional shares of common stock.

Incentive stock options to purchase 40 million shares of common stock after July 1, 2020, at $14 per share, were outstanding at the beginning and end of 2021. The average market price of Canadays common stock was $ 20 per share during 2021.

Canaday's net income for the year ended December 31, 2021, was $1500 million, the effective income tax rate was 25%.
Answer #1

Calculation of Basic EPS

Numerator (Basic EPS): Net income = $1500 million;

Preferred dividends = $60 million [(6% x $50) x 20 million]. Because the preferred stock is cumulative dividends are included whether or not paid.

Denominator (Basic EPS): Weighted average # shares common stock outstanding

1/1 – 12/31 660 million x (12/12) = 660

9/1 – 12/31 72 million x (4/12) = 24

Weighted average # shares 684

Basic EPS = ($1500 - $60) ÷ 684 = $2.105

Calculation of Diluted EPS

Stock Options: Use the Treasury Stock Method if dilutive.

Step 1. Are the stock options dilutive? Yes because the exercise price of $14 is less than the average market price of $20.

Step 2. Assume exercise at the later of the date of issue (7/1/20) or the beginning of the period (1/1/21). Assume exercise 1/1/21.

Step 3. Proceeds received on exercise = 40 million x $14 = $560 million

Step 4. # Shares repurchased = $560 million ÷ $20 = 28 million

Step 5. Net increase in # shares outstanding = 12 million (40 million – 28 million)

Step 6. EPS with the inclusion of the options:

($1,500 - $60) ÷ (684 + 12) = $2.068 (this will be used to test for dilutive effect of convertible bonds).

Convertible Bonds: Use the If Converted Method

Step 1. Assume conversion at later of date of issue (?) or beginning of year (1/1/21). Assume conversion on 1/1/21.

Step 2. Interest not paid net of tax = $168 million [(6.4% x $3500 million) x 75%]

Step 3. additional shares on conversion = 70 million

Step 4. Conversion ratio = $168 ÷ 70 = $2.40

Step 5. Dilutive Stock Option because $2.068 is less than $2.40

Diluted EPS=(Net Income−Preferred Dividends) /(Weighted Average Shares Outstanding+Conversion of dilutive securities)

(1500-60) /(684+70)=1.909

Basic EPS = ($1500 - $60) ÷ 684 = $2.105

Diluted EPS = (1500-60) /(684+70) = 1.909

Ask any query and Please give Thumbs up